OpenStudy (helpblahblahblah):

ILL GIVE MEDAL: What is 36% of 50? A. 1.4 B. 0.72 C. 18 D. 72

OpenStudy (helpblahblahblah):

i got 18 is that correct?

OpenStudy (helpblahblahblah):

are you going to help @beastmachine ?

OpenStudy (anonymous):

i believe so because that is what i just got also and srry for the wait i was doing something else at the moment

OpenStudy (helpblahblahblah):

its okay and thanks :)

OpenStudy (doc.brown):

36 is to 100 as what is to 50? 36/100 = x/50

OpenStudy (doc.brown):

or you have 36 100ths of 50 x = 50 * 36/100

OpenStudy (doc.brown):

or chop 50 into 100 pieces and only take 36 of those pieces. x = 50/100 *36

OpenStudy (helpblahblahblah):

wait what @doc.brown

OpenStudy (anonymous):

what are you talking about @doc.brown all you have to do is divide 50 by 36........... and that is your answer

OpenStudy (doc.brown):

Chopping 50 into 36 pieces gives you 1/36 of 50. We want 36% of 50.